OpenStudy (solomonzelman):

Why is the natural log function "ln" is absent in wolfram? it always uses "log"....

OpenStudy (solomonzelman):

Like for example: http://www.wolframalpha.com/input/?i=ln%28x%29

OpenStudy (solomonzelman):

or is it meant to be just \(\large\color{teal}{ \log_e }\) ?

OpenStudy (turingtest):

if you use log it defaults to natural log

OpenStudy (anonymous):

well i guess they prefer using logs instead of ln.. beside dont you get the same ourput?

OpenStudy (solomonzelman):

ln(x) is not the same thing as log(x), you know that better than me....

OpenStudy (solomonzelman):

TuringTest's reasoning seems to be legit:)

OpenStudy (turingtest):

some people use ln so much they just call it log, and I guess mr wolfram is one of those people lol if you want to control the base type log_2(2)

OpenStudy (turingtest):

or log_x(y)

OpenStudy (solomonzelman):

I know how to control tnx base, tnx;) if you need a question about how to put integrals sigmas and anything inline, I can assist:P

OpenStudy (solomonzelman):

people mostly think that you can't do this: \(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ f(n)}\)

OpenStudy (turingtest):

oh thanks, I'm sure i'll need help from time to time im gonna guess it is similar to latex, but i'm sure id have to play with it

OpenStudy (solomonzelman):

`\(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ f(n)}\)` is the sigma code

OpenStudy (turingtest):

for latex you mean

OpenStudy (solomonzelman):

Yes, turing;) @Marki, you need this, use it. Inline sigma, very good for you

OpenStudy (turingtest):

sum j^2, j=1 to 1000 would work on wolfy :)

OpenStudy (solomonzelman):

yes, but to get a better vision and think good of yourself, ``` \(\large\color{black}{\displaystyle\lim_{x \rightarrow ~a}f(x)}\) LIMIT \(\large\color{black}{\displaystyle\int\limits_{~}^{~}~~~dx}\) INTEGRAL \(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ f(n)}\) SIGMA ```

OpenStudy (anonymous):

thanks solmon :) the thing is im too lazy to write :P

OpenStudy (solomonzelman):

(they are all completely blank, if that is inconvenient, sorry)

OpenStudy (solomonzelman):

anyway, this is not meant to be a latex tutorial;) tnx for viewing....

OpenStudy (solomonzelman):

any question now are available though here or in messages.

OpenStudy (solomonzelman):

I can make an attempt to put a cube

OpenStudy (anonymous):

ok here is one thing the default log in wolfram is ln :) (base e not 10)

OpenStudy (solomonzelman):

lost connection and the cube

OpenStudy (solomonzelman):

oh, there is no simple trick: doesn't work with \angle{} \text{ ___} but I can make a square though \(\normalsize\color{ slate }{\Huge{\bbox[5pt, lightcyan ,border:2px solid black ]{ \color{lightcyan}{\LARGE\frac{~~\frac{~}{~~~\frac{~\frac{}{}~}{~\frac{~}{~}}~~~~}~~~}{~\frac{\frac{~}{~\frac{~\frac{}{~}}{~}}~~}{~}~~} } }}}\)

OpenStudy (solomonzelman):

okay, done

OpenStudy (solomonzelman):

yes, marki base e, what I asked to be an option at the beginnn, tnx

OpenStudy (anonymous):

\(\normalsize\color{ slate }{\Huge{\bbox[5pt, lightcyan ,border:2px solid black ]{ \color{lightcyan}{\LARGE\frac{~~\frac{~}{~~~\frac{~\frac{}{}~}{~\frac{~}{~}}~~~~}~~~}{~\frac{\frac{~}{~\frac{~\frac{}{~}}{~}}~~}{~}~~} } }}} \) NICE!